OpenStudy (anonymous):

The 1st, 5th, and 13th terms of an arithmetic sequence are the first three terms of a geometric sequence with a common ratio of 2. If the 21st term of the arithmetic sequence is 72, calculate the sum of the first 10 terms of the geometric sequence.

OpenStudy (here_to_help15):

the 1st term of the A.S. is a, and the 5th term is a + 4k from the G.S. with a common ratio of 2 (a + 4k)/a = 2 a + 4k = 2a a = 4k then from the 21st term of the A.S. a + 20k = 72 substitute a = 4k 4k + 20k = 72 24k = 72 k = 3 a = 4*3 a = 12 the sum of the first 10 terms of the G.S. is S_10 = 12*(1 - 2^10)/(1 - 2) S_10 = 12276